# Timelike entanglement entropy with gravitational anomalies

### v1

@@ 3.2 Finite temperature and angular potential
We find that the above TEE is complex where only one central charge contributes to the imaginary part and both central charges are present in the real part, similar to the zero temperature case. However the anomaly contribution to the real part (proportional to $c_{L}-c_{R}$) is present in contrast to earlier case. It matches with the CFT2 result obtained in [23] when the anomaly is absent i.e. $c_{L}=c_{R}$. We note that in the zero temperature limit $(\beta_{L},\beta_{R})\gg T$, the above expression reduces to (3.5). We also note that one can consider the Rindler approach [43] in the presence of gravitational anomaly and obtain the same results (3.7) and (3.5).
@@ 4 Holographic Timelike entanglement entropy in AdS3/CFT2 with gravitational anomaly
In this section, we compute the timelike entanglement entropy holographically in the context of the AdS/CFT correspondence for dual conformal field theories with a gravitational anomaly. In this case the bulk dual geometry is described by topologically massive gravity (TMG) in a bulk AdS3 spacetime [12, 4].
@@ 4.1 Review of earlier works
We first briefly review about the holographic description in the context of Topologically Massive Gravity (TMG) in AdS3. The bulk action for TMG in AdS3 consists of the standard Einstein-Hilbert action and an additional gravitational Chern-Simons (CS) term which is given by [68, 69, 4]
@@
where $C_{\mu\nu}$ denotes the Cotton tensor. When the Cotton tensor vanishes, the theory admits solutions that retain the form of Einstein metrics ensuring that these solutions remain locally AdS3. In this work, we confine our discussion to such locally AdS3 solutions. The asymptotic symmetry analysis of TMG in AdS3 reveals that the algebra of asymptotic Killing vector modes corresponds to two copies of the Virasoro algebra with left and right moving central charges. This results in a Brown-Henneaux [70] type relation for TMG-AdS3 given by [71, 72]
@@
For AdS/CFT correspondence, the authors [4] proposed a bulk description for the holographic entanglement entropy in the presence of gravitational anomalies by extending the Lewkowycz-Maldacena prescription [73]. They showed that the holographic entanglement entropy in such cases can be obtained by extremizing a worldline action for a spinning particle propagating in the AdS3 bulk. Consequently, the holographic entanglement entropy is given by the on-shell action of a massive spinning particle in the topologically massive gravity as
[…]
where $C$ represents the worldline of the particle. Here $n_{1}$ and $n_{2}$ are the normal vectors to the curve (up to a choice of handedness) such that $n_{1}\cdot n_{2}=0$. The subscript $E$ indicates that the above expression is evaluated in the Euclidean signature. For locally AdS3 spacetimes where the worldline that minimizes the action is a geodesic, the first part in the above equation corresponds to the Ryu-Takayanagi term originating from the Einstein-Hilbert action and the second part arises from the Chern-Simons term.
When the path (curve) $C$ is spacelike, one can choose $n_{1}=\partial_{t}$ and $n_{2}=\partial_{x}$ where they are both spacelike in the Euclidean signature. On analytically continuing to the complex coordinates $(z,\bar{z})$ via $z=x-t,\bar{z}=x+t$, the normal vectors $(n,\tilde{n})$ in Lorentzian signature is given by
@@
In Lorentzian signature, one of the normal vector becomes timelike $(n)$ and the other remains spacelike $(\tilde{n})$. The entanglement entropy (4.4) in terms of these normal vectors can be written as
[…]
@@
where $n_{i}$ and $n_{f}$ are normal vectors at the end point of the path (or geodesic). As described in [4], the integral (4.7) measures the twist of $n_{f}$ relative to $n_{i}$ along $C$. To obtain a compact expression for this integral, one can introduce a vector $n(s)$ that obeys the boundary condition (4.8) and a normal frame along the path defined by the pair of vectors $(q^{\mu},\tilde{q}^{\mu})$ which are both parallel transported i.e. $\nabla q=\nabla\tilde{q}=0$. The normal vector $n(s)$ can then be written in terms of $q$ and $\tilde{q}$ as
@@
where $\eta$ is the rapidity of the Lorentz boost required to boost the orthonormal frame $(v,q,\tilde{q})$ to $(v,n,\tilde{n})$. The integral (4.7) then simplifies to a total derivative given by
@@
The anomalous contribution from the spacelike curve can be written in terms of $q$ and $\tilde{q}$ by utilizing (4.9) and (4.10) as [4]
@@
The holographic entanglement entropy is determined only by spacelike geodesic in the bulk and hence, it is obtained by extremizing the on-shell action (4.6) as follows [4]
@@ 4.2 Anomalous contribution from timelike geodesic
Having described the anomalous contribution from the spacelike geodesic to the holographic entanglement entropy, we now proceed to derive the anomalous (or CS) contribution from timelike geodesic to the on-shell action (4.4). For a timelike path (or geodesic) denoted as $C_{\tau}$, the tangent vector is timelike and the other two normal vectors are spacelike which is different from the spacelike geodesic case. We now choose the normal vectors $(n,\tilde{n})$ in (4.4) as
@@
These vectors remain spacelike under analytic continuation to Lorentzian signature. We now find the following expression for the on-shell action for a timelike curve in terms of these vectors in the Lorentzian signature using (4.4) as
[…]
@@
[…]
Similar to the earlier section, the integration only depend on the boundary terms i.e. it depends only on the the initial and the final values of the normal vector $n$ (or $\tilde{n}$). We set the boundary values of the normal vector $n(\tau)$ as
@@
where $n_{i}$ and $n_{f}$ are normal vectors located at the end point of the curve. As both the normal vectors ($n,\tilde{n}$) are spacelike for this case, they rotate along the curve as we parallel transport them from initial to the final point. In other words when we parallel transport $n_{i}$ to the final point, the resulting vector is related to $n_{f}$ via $SO(2)$ transformation. [^3] They are related via $SO(1,1)$ transformation for the case of spacelike curves where one normal vector is timelike and other is spacelike [4]. Now proceeding similar to spacelike curve case, consider a vector $n(\tau)$ that satisfies the boundary condition (4.17) and a parallel transported normal frame which consists of two vectors $(q^{\mu},\tilde{q}^{\mu})$ satisfying $\nabla q=\nabla\tilde{q}=0$. We can expand $n(\tau)$ in terms of $q$ and $\tilde{q}$ as
@@
[…]
@@
$v^{2}=-1$ $,\quad q^{2}=\tilde{q}=1,\quad q\cdot\tilde{q}=v\cdot q=v\cdot\tilde{q}=0$ (4.21)
@@ 4.3 Holographic TEE at zero temperature
where $\ell$ is the AdS radius. As discussed in [28] for the case of AdS3/CFT2, the geodesics required for computing the holographic TEE consists of two spacelike geodesics connecting the endpoints of $A$ and infinities plus a timelike geodesic which connects the endpoints of two spacelike geodesics. The equation for these geodesics in the Poincaré metric (4.23) is given by
@@
where $T$ and $R$ are constants. The curve (4.24a) is spacelike and describes a geodesics that starts from the end point $t=\pm T/2$ and ends at infinity. These spacelike geodesics are smoothly joined at infinity by a timelike geodesic (4.24b) for any $R>0$. It is depicted in fig. 2.
@@
The equation for the spacelike geodesics as denoted by red color in fig. 2 is given by (4.24a). The total length of these geodesics can be obtained as
@@
These vectors satisfy the condition given in (4.12). While the normal vector $n$ can be readily fixed at the boundary along the time direction as described in [4], doing so at the geodesic’s other end in the bulk is not straightforward because of the non-covariance of bulk TMG, requiring a particular prescription. The normal vectors at the endpoints of the relevant geodesics is determined as follows. First, we choose an orthonormal triad $(v^{\mu},n^{\mu},\tilde{n}^{\mu})$ at the initial endpoint of interval where $v$ represents the tangent vector, $n$ is aligned with the boundary future time direction and $\tilde{n}$ is determined by orthogonality and the right hand convention. For the other endpoint of spacelike geodesic I as depicted in fig. 2, the vector $\tilde{n}$ is maintained as perpendicular to the plane (along $x$ direction) and $n$ is subsequently determined based on the tangent vector $v$ and $\tilde{n}$. Secondly, at the final endpoint of the interval, an orthonormal triad is established with $n$ remaining along the future time direction and $\tilde{n}$ derived from $n$ and $v$ which is now oriented towards the negative x direction. This specific $\tilde{n}$ is then utilized at the other endpoint of spacelike geodesic II to determine the normal vector $n$ using $v$ and $\tilde{n}$. Finally, these determined normal vectors at the endpoints of the spacelike geodesics in the bulk are smoothly connected across null infinity to find the normal vectors $n$ at the endpoints of the timelike geodesic.
For the spacelike geodesic I as shown in fig. 2 which starts from $t=-T/2$ at the boundary ($u\to 0$) and goes to infinity $u\to\infty$, it can be parameterised by $\beta$ as $t=-T/2\cosh\beta$ and $u=T/2\sinh\beta$ such that $\beta=0$ corresponds to the boundary and $\beta\to\infty$ for the other end point of the geodesic. The normal vector $n$ is timelike for the case of spacelike geodesic as required by condition (4.5). Since the other end point is at infinity and normal vector becomes null at $\beta\to\infty$, we regularize it by taking some finite but large value of $\beta$ so that it remains timelike near infinity and then take $\beta\to\infty$ at the end of computations. The boundary values of the (normalised) normal vector $n$ can be written as
@@
where the initial value of normal vector at boundary is same as considered in [4] and determined by the boundary CFT data. The final value of $n$ is perpendicular to geodesic at null infinity (as the geodesic approaches the line $t=u$ at $u\to\infty$).
@@
This is very similar to the case for the entanglement entropy for an unboosted interval (i.e a pure spacelike interval) [4] where anomalous contribution also vanishes. Therefore, the contribution from the spacelike geodesics to the on-shell action (4.6) is given by
@@
For the timelike geodesic (green color curve in fig. 3) as described by (4.24b), the length of this geodesic can be obtained as
@@
The timelike geodesic (4.24b) can be parameterized by $\gamma$ as $u=R\cosh\gamma,t=R\sinh\gamma$ where $\gamma$ range from $\infty$- ∞ to +$\infty$+ ∞. Given that the timelike geodesic matches up with the spacelike geodesics I, II asymptotically at the null infinities, we propose to specify the initial and final normal vectors $n_{1}$, $n_{2}$ by matching them up with the normal vectors $n_{i}$ and $n_{f}$ at the respective null infinities (point 1 and 2 in fig. 3). As such we have
@@
where $\gamma$ goes to $\infty$- ∞ near the initial point 1. As the endpoints extend to infinity, we introduce a regularization for computational purposes by choosing $\gamma$ a large value similar to the previous subsection. This ensures the boundary value of normal vectors remain spacelike near infinity i.e at the intersection of spacelike and timelike geodesics. Similarly we have
@@
$S_{\rm anom}^{\rm timelike}$ $=\frac{1}{4G_{N}^{(3)}\mu}\log\left[\frac{q(s_{2})\cdot n_{2}-i\tilde{q}(s_{2})\cdot n_{2}}{q(s_{1})\cdot n_{1}-i\tilde{q}(s_{1})\cdot n_{1}}\right]$ (4.35)

### v2

@@ 3.2 Finite temperature and angular potential
We find that the above TEE is complex where only one central charge contributes to the imaginary part and both central charges are present in the real part, similar to the zero temperature case. However the anomaly contribution to the real part (proportional to $c_{L}-c_{R}$) is present in contrast to earlier case. It matches with the CFT2 result obtained in Doi:2023zaf when the anomaly is absent i.e. $c_{L}=c_{R}$. We note that in the zero temperature limit $(\beta_{L},\beta_{R})\gg T$, the above expression reduces to (3.5). We also note that one can consider the Rindler approach He:2023ubi in the presence of gravitational anomaly and obtain the same results (3.7) and (3.5).
@@ 4 Holographic Timelike entanglement entropy in AdS3/CFT2 with gravitational anomaly
In this section, we compute the timelike entanglement entropy holographically in the context of the AdS/CFT correspondence for dual conformal field theories with a gravitational anomaly. In this case the bulk dual geometry is described by topologically massive gravity (TMG) in a bulk AdS3 spacetime Kraus:2005zm ; Castro:2014tta .
@@ 4.1 Review of earlier works
We first briefly review about the holographic description in the context of Topologically Massive Gravity (TMG) in AdS3. The bulk action for TMG in AdS3 consists of the standard Einstein-Hilbert action and an additional gravitational Chern-Simons (CS) term which is given by Tachikawa:2006sz ; Skenderis:2009nt ; Castro:2014tta
@@
where $C_{\mu\nu}$ denotes the Cotton tensor. When the Cotton tensor vanishes, the theory admits solutions that retain the form of Einstein metrics ensuring that these solutions remain locally AdS3. In this work, we confine our discussion to such locally AdS3 solutions. The asymptotic symmetry analysis of TMG in AdS3 reveals that the algebra of asymptotic Killing vector modes corresponds to two copies of the Virasoro algebra with left and right moving central charges. This results in a Brown-Henneaux Brown:1986nw type relation for TMG-AdS3 given by Hotta:2008yq ; Compere:2008cv
@@
For AdS/CFT correspondence, the authors Castro:2014tta proposed a bulk description for the holographic entanglement entropy in the presence of gravitational anomalies by extending the Lewkowycz-Maldacena prescription Lewkowycz:2013nqa . They showed that the holographic entanglement entropy in such cases can be obtained by extremizing a worldline action for a spinning particle propagating in the AdS3 bulk. Consequently, the holographic entanglement entropy is given by the on-shell action of a massive spinning particle in the topologically massive gravity as
$S_{\rm on-shell}=\frac{1}{4G_{N}^{(3)}}\int_{C}ds\left(\sqrt{g_{\mu\nu}\dot{X}^{\mu}\dot{X}^{\nu}}+\frac{i}{\mu}n_{2}\cdot\nabla n_{1}\right)_{E},$ (4.4)
where $C$ represents the worldline of the particle. Here $n_{1}$ and $n_{2}$ are the normal vectors to the curve (up to a choice of handedness) such that $n_{1}\cdot n_{2}=0$ and $\nabla$ denotes the covariant derivative along the worldline which is defined as
$\nabla V^{\mu}\equiv\frac{dV^{\mu}}{ds}+\Gamma^{\mu}_{\lambda\rho}\frac{dX^{\rho}}{ds}V^{\lambda}.$ (4.5)
The subscript $E$ indicates that the above expression is evaluated in the Euclidean signature. For locally AdS3 spacetimes where the worldline that minimizes the action is a geodesic, the first part in the above equation corresponds to the Ryu-Takayanagi term originating from the Einstein-Hilbert action and the second part arises from the Chern-Simons term.
When the path (curve) $C$ is spacelike, one can choose $n_{1}=\partial_{t}$ and $n_{2}=\partial_{x}$ where they are both spacelike in the Euclidean signature. On analytically continuing to the complex coordinates $(z,\bar{z})$ via $z=x-t,\bar{z}=x+t$, the normal vectors (n,n~)($\tilde{n}$) in Lorentzian signature is given by
@@
In Lorentzian signature, one of the normal vector becomes timelike $(n)$ and the other remains spacelike (n~)($\tilde${n}). The entanglement entropy (4.4) in terms of these normal vectors can be written as
$S_{\rm on-shell}^{\rm spacelike}=\frac{1}{4G_{N}^{(3)}}\int_{C}ds\left(\sqrt{g_{\mu\nu}\dot{X}^{\mu}\dot{X}^{\nu}}+\frac{1}{\mu}\tilde{n}\cdot\nabla n\right).$ (4.7)
@@
where $n_{i}$ and $n_{f}$ are normal vectors at the end point of the path (or geodesic). As described in Castro:2014tta , the integral (4.8) measures the twist of $n_{f}$ relative to $n_{i}$ along $C$. To obtain a compact expression for this integral, one can introduce a vector $n(s)$ that obeys the boundary condition (4.9) and a normal frame along the path defined by the pair of vectors $(q^{\mu},\tilde{q}^{\mu})$ which are both parallel transported i.e. $\nabla q=\nabla\tilde{q}=0$. The normal vector $n(s)$ can then be written in terms of $q$ and $\tilde{q}$ as
@@
$S_{\rm anom}^{\rm spacelike}=\frac{1}{4G_{N}^{(3)}\mu}\int ds\,\dot{\eta}(s)=\frac{1}{4G_{N}^{(3)}\mu}\left(\eta(s_{f})-\eta(s_{i})\right).$ (4.11)
The anomalous contribution from the spacelike curve can be written in terms of $q$ and $\tilde{q}$ by utilizing (4.10) and (4.11) as Castro:2014tta
@@
The holographic entanglement entropy is determined only by spacelike geodesic in the bulk and hence, it is obtained by extremizing the on-shell action (4.7) as follows Castro:2014tta
@@ 4.2 Anomalous contribution from timelike geodesic
Having described the anomalous contribution from the spacelike geodesic to the holographic entanglement entropy, we now proceed to derive the anomalous (or CS) contribution from timelike geodesic to the on-shell action (4.4). For a timelike path (or geodesic) denoted as $C_{\tau}$, the tangent vector is timelike and the other two normal vectors are spacelike which is different from the spacelike geodesic case. We now choose the normal vectors (n,n~)($\tilde{n}$) in (4.4) as
@@
These vectors remain spacelike under analytic continuation to the Lorentzian signature. We now find the following expression for the on-shell action for a timelike curve in terms of these vectors in the Lorentzian signature using (4.4) as
$S_{\rm on-shell}^{\rm timelike}=\frac{i}{4G_{N}^{(3)}}\int_{C_{\tau}}d\tau\left(\sqrt{g_{\mu\nu}\dot{X}^{\mu}\dot{X}^{\nu}}+\frac{1}{\mu}\tilde{n}\cdot\nabla n\right),$ (4.16)
@@
$S_{\rm anom}^{\rm timelike}=\frac{i}{4G_{N}^{(3)}\mu}\int_{C_{\tau}}d\tau\;\epsilon_{\mu\nu\rho}v^{\mu}n^{\nu}(\nabla n^{\rho}).$ (4.17)
Similar to the earlier section, the integration only depend on the boundary terms i.e. it depends only on the initial and the final values of the normal vector $n$ (or $\tilde{n}$). We set the boundary values of the normal vector $n(\tau)$ as
@@
where $n_{i}$ and $n_{f}$ are normal vectors located at the end point of the curve. As both the normal vectors ($n,\tilde{n}$) are spacelike for this case, they rotate along the curve as we parallel transport them from initial to the final point. In other words when we parallel transport $n_{i}$ to the final point, the resulting vector is related to $n_{f}$ via $SO(2)$ transformation. [^3] They are related via $SO(1,1)$ transformation for the case of spacelike curves where one normal vector is timelike and other is spacelike Castro:2014tta . Now proceeding similar to spacelike curve case, consider a vector $n(\tau)$ that satisfies the boundary condition (4.18) and a parallel transported normal frame which consists of two vectors $(q^{\mu},\tilde{q}^{\mu})$ satisfying $\nabla q=\nabla\tilde{q}=0$. We can expand $n(\tau)$ in terms of $q$ and $\tilde{q}$ as
@@
$S_{\rm anom}^{\rm timelike}=\frac{i}{4G_{N}^{(3)}\mu}\int d\tau\,\dot{\theta}(\tau)=\frac{i}{4G_{N}^{(3)}\mu}\left(\theta(\tau_{f})-\theta(\tau_{i})\right).$ (4.20)
@@
$v^{2}=-1$ $,\quad q^{2}=\tilde{q}^{2}=1,\quad q\cdot\tilde{q}=v\cdot q=v\cdot\tilde{q}=0$ (4.22)
@@ 4.3 Holographic TEE at zero temperature
where $\ell$ is the AdS radius. As discussed in Basak:2023otu for the case of AdS3/CFT2, the geodesics required for computing the holographic TEE consists of two spacelike geodesics connecting the endpoints of $A$ and infinities plus a timelike geodesic which connects the endpoints of two spacelike geodesics. The equation for these geodesics in the Poincaré metric (4.24) is given by
@@
where $T$ and $R$ are constants. The curve (4.25a) is spacelike and describes a geodesics that starts from the end point $t=\pm T/2$ and ends at infinity. These spacelike geodesics are smoothly joined at infinity by a timelike geodesic (4.25b) for any $R>0$. It is depicted in fig.˜2.
@@
The equation for the spacelike geodesics as denoted by red color in fig.˜2 is given by (4.25a). The total length of these geodesics can be obtained as
@@
These vectors satisfy the condition given in (4.13). While the normal vector $n$ can be readily fixed at the boundary along the time direction as described in Castro:2014tta , doing so at the geodesic’s other end in the bulk is not straightforward because of the non-covariance of bulk TMG, requiring a particular prescription. The normal vectors at the endpoints of the relevant geodesics is determined as follows. First, we choose an orthonormal triad $(v^{\mu},n^{\mu},\tilde{n}^{\mu})$ at the initial endpoint of interval where $v$ represents the tangent vector, $n$ is aligned with the boundary future time direction and $\tilde{n}$ is determined by orthogonality and the right hand convention. For the other endpoint of spacelike geodesic I as depicted in fig.˜2, the vector $\tilde{n}$ is maintained as perpendicular to the plane (along $x$ direction) and $n$ is subsequently determined based on the tangent vector $v$ and $\tilde{n}$. Secondly, at the final endpoint of the interval, an orthonormal triad is established with $n$ remaining along the future time direction and $\tilde{n}$ derived from $n$ and $v$ which is now oriented towards the negative x direction. This specific $\tilde{n}$ is then utilized at the other endpoint of spacelike geodesic II to determine the normal vector $n$ using $v$ and $\tilde{n}$. Finally, these determined normal vectors at the endpoints of the spacelike geodesics in the bulk are smoothly connected across null infinity to find the normal vectors $n$ at the endpoints of the timelike geodesic.
For the spacelike geodesic I as shown in fig.˜2 which starts from t=T/2t=-T/2 at the boundary ($u\to 0$) and goes to infinity $u\to\infty$, it can be parameterised by $\beta$ as $t=-T/2\cosh\beta$ and $u=T/2\sinh\beta$ such that $\beta=0$ corresponds to the boundary and $\beta\to\infty$ for the other end point of the geodesic. The normal vector $n$ is timelike for the case of spacelike geodesic as required by condition (4.6). Since the other end point is at infinity and normal vector becomes null at $\beta\to\infty$, we regularize it by taking some finite but large value of $\beta$ so that it remains timelike near infinity and then take $\beta\to\infty$ at the end of computations. The boundary values of the (normalised) normal vector $n$ can be written as
@@
where the initial value of normal vector at boundary is same as considered in Castro:2014tta and determined by the boundary CFT data. The final value of $n$ is perpendicular to geodesic at null infinity (as the geodesic approaches the line t=ut=-u at $u\to\infty$).
@@
This is very similar to the case for the entanglement entropy for an unboosted interval (i.e a pure spacelike interval) Castro:2014tta where anomalous contribution also vanishes. Therefore, the contribution from the spacelike geodesics to the on-shell action (4.7) is given by
@@
For the timelike geodesic (green color curve in fig.˜3) as described by (4.25b), the length of this geodesic can be obtained as
@@
The timelike geodesic (4.25b) can be parameterized by $\gamma$ as $u=R\cosh\gamma,t=R\sinh\gamma$ where $\gamma$ range from -$\infty$ to $\infty$. Given that the timelike geodesic matches up with the spacelike geodesics I, II asymptotically at the null infinities, we propose to specify the initial and final normal vectors $n_{1}$, $n_{2}$ by matching them up with the normal vectors $n_{i}$ and $n_{f}$ at the respective null infinities (point 1 and 2 in fig.˜3). As such we have
@@
where $\gamma$ goes to -$\infty$ near the initial point 1. As the endpoints extend to infinity, we introduce a regularization for computational purposes by choosing $\gamma$ a large value similar to the previous subsection. This ensures the boundary value of normal vectors remain spacelike near infinity i.e at the intersection of spacelike and timelike geodesics. Similarly we have
@@
$S_{\rm anom}^{\rm timelike}$ $=\frac{1}{4G_{N}^{(3)}\mu}\log\left[\frac{q(\tau_{2})\cdot n_{2}-i\tilde{q}(\tau_{2})\cdot n_{2}}{q(\tau_{1})\cdot n_{1}-i\tilde{q}(\tau_{1})\cdot n_{1}}\right]$ (4.36)